OpenStudy (anonymous):

PLEASE HELP and ill MEDAL!!Simplify the imaginary unit i. Square root of -16

OpenStudy (shiraz14):

4i

OpenStudy (anonymous):

\[\sqrt{-16}=\sqrt{16*-1}=\sqrt{16}*\sqrt{-1}=4*-1=4i\]

OpenStudy (anonymous):

oh thats all u do?

OpenStudy (anonymous):

yeah sqrt of -1=i

OpenStudy (shiraz14):

That's correct @xlovely_Lizardx

OpenStudy (anonymous):

oh yeah i remember that! i wrote that in my notes

OpenStudy (anonymous):

Mind helping me with one more????

OpenStudy (anonymous):

sure :D

OpenStudy (radar):

\[\sqrt{16}\sqrt{-1}=4\sqrt{-1}=4i\]

OpenStudy (anonymous):

-5+i over 2i ...simplify the expression.

OpenStudy (anonymous):

hmmm I havn't done imaginary stuff in a few years but if i remember correctly:\[\frac{ -5+i }{ 2i }=\frac{ -5+\sqrt{-1} }{ \sqrt{-2} }=-5+\sqrt{\frac{ -1 }{ -2 }}=-5+\sqrt{\frac{ 1 }{ 2 }}\]

OpenStudy (anonymous):

does that look right?

OpenStudy (anonymous):

i think..lemme check my work

OpenStudy (anonymous):

i have noo idea. @radar do you kknow if this is right?????????

OpenStudy (anonymous):

@radar

OpenStudy (radar):

I'm here (now). Multiply the numerator and denominator by i i/i if 1 and would not change the value of your fraction. this would give you (-5i + i^2)/2i^2 Do you follow ?

OpenStudy (anonymous):

yes i have taht

OpenStudy (radar):